# Performance of Quantum Approximate Optimization with Quantum Error Detection

### v1

@@ III.2.2 Fitted error rates
The deviations between the fitted error rates and the emulator noise rates presented in Table 1 contain valuable information about where other noise sources accumulate.
Starting from the error rate $p_{cx}$ of cnot gates, the fitted value is more than four times larger than the emulator error rate, showing that a significant amount of noise unaccounted by the performance model accumulates in the error detection blocks of the Iceberg code. For the logical gates we consider two noise channels, with error rates $p_{c}$ and $p_{a}$, that introduce Pauli errors that commute and anti-commute, respectively, with the two Iceberg code stabilizers. Since the values established in the emulator depend on the rotation angles of the logical gates and the QAOA circuits do not present a clear tendency towards any particular angle, Table 1 presents the minimum and maximum values. We find that these fitted error rates are almost twice larger than the maximum value given by the emulator, hinting again that the logical gates in the Iceberg code circuit accumulate unnacounted noise. In contrast, the fitted value of the error rate $p_{\ell}$ of the unencoded logical gates is well approximated by the minimum value established by the emulator.
@@
The values reported for the emulator in Table 1 are obtained from the parameters of the depolarising channel for the native two-qubit gates $\exp(-i\theta ZZ)$ and their dependence with the QAOA rotation angles $\theta\in\{\gamma,\beta\}$. The depolarising channel assigns a probability $q_{\sigma}$ to each of the 15 Pauli errors in $\sigma=\mathcal{P}^{\otimes 2}\setminus\{I^{\otimes 2}\}$ after the native gate. The total probability of a Pauli error is the value of $p_{cx}$ reported in the table. These values are corrected by a multiplicative factor defined as linearly increasing function $r(\theta)\simeq a+b|\theta|$ in the angle magnitude, such that $r(\pi/4)=1$ for maximally entangling gates like the cnot. The error rate of unencoded logical gates is then $p_{\ell}(\theta)=p_{cx}r(\theta)$. For the Iceberg code we additionally separate the error rate of commuting errors $q_{c}=q_{X^{\otimes 2}}+q_{Y^{\otimes 2}}+q_{Z^{\otimes 2}}$ from that of anti-commuting errors $q_{a}=p_{cx}-q_{c}$. To unify with the model error rates, we factorise this single channel into a product of a commuting channel with error rate $p_{c}(\theta)=q_{c}r(\theta)(1-q_{a}r(\theta))$ and an anti-commuting channel with error rate $p_{a}(\theta)=q_{a}r(\theta)$. We report the minimum and maximum among all QAOA rotation angles.
@@
As observed from the shift of the breakeven frontiers (red lines) in Fig. 3A, the QAOA logical fidelity stabilizes as the number of syndrome measurements increases, even though the post-selection rate, indicated by the orange lines in Fig. 3C, decreases. This aligns with our experimental data and findings in the literature [18], which suggest that the initial syndrome measurements significantly enhance circuit performance, while the marginal gains diminish with an increasing number of syndrome measurements.
Figure 4: Hardware improvement necessary for QAOA to become competitive with the Goemans-Williamson algorithm. A Solve $k=16$ MaxCut using different solvers. Each data is reported as the mean of approximation ratio over 100 $k=16$ $3$-regular graphs. The standard errors are too small to be seen. B Scaling of model parameter to beat Goemans-Williamson (GW) algorithm for $k=16$ graphs. The Iceberg code helps the QAOA to beat GW earlier than an unencoded one.
@@ III.2.4 Predicting performance on future hardware
As shown in Fig. 3B and D, as the factor decreases, we observe a significant shift of the performance frontier to larger number $k$ of logical qubits while the post-selection rate improves dramatically. This indicates that with higher-quality quantum hardware, we can push the breakeven frontier of logical fidelity to deeper circuits on larger problem instances with less post-selection overhead.
To elucidate the conditions for QAOA to become competitive with classical solvers, we use our model to answer the question of when a QAOA hardware experiment can outperform the Goemans-Williamson (GW) algorithm [57] in terms of approximation ratio. As reported in the literature [59], a noiseless QAOA with fixed parameters has been able to surpass the GW algorithm on small graphs. In Fig. 4A, we show an example of solving $k=16$-node 3-regular graphs with noiseless QAOA, GW, and Iceberg QAOA with four syndrome measurements as well as unencoded QAOA emulated on the H2-1 emulator. The noiseless QAOA is able to outperform GW for $\ell\geq 6$ layers. However, both the Iceberg and unencoded QAOA have not yet surpassed GW.
Specifically, at $\ell=10$, the approximation ratio of noiseless QAOA is $0.9810\ldots$, while the average approximation ratio for GW is $0.9554\ldots$. This implies that the logical fidelity of a noisy QAOA must be of at least $\nicefrac{{0.9554}}{{0.9810}}\simeq 0.974$ to outperform GW, assuming our white noise model approximation. In Fig. 4B, we vary the scaling factor of the model parameters to determine when this breakeven logical fidelity can be achieved. The results indicate that the Iceberg and unencoded circuits require scaling factors of approximately $0.81$ and $0.60$, respectively. This suggests that the Iceberg code enables a breakeven fidelity on small MaxCut problems much earlier than unencoded circuits as hardware technology advances.
Figure 5: Comparison between experimental data and model predictions on random Erdős–Rényi graphs with different numbers of edges. The prediction of Iceberg logical fidelity is less accurate compared to testing on 3-regular graphs, while the prediction of unencoded logical fidelity and the prediction of post-selection rate remain accurate.
@@ III.2.5 Model generalization beyond $3$-regular graphs
We present the comparison between experimental results and model predictions in Fig. 5. The model predictions for the logical fidelity of the Iceberg code circuits are less accurate, specially with the densest graph of $|E|=45$ edges. The model predictions for the unencoded logical fidelity and the Iceberg code post-selection rates are comparably more accurate. This indicates that the fitted model works well for problems with similar topologies, but highlights the limitation of the model’s generalization to different problems. We suspect that the worse model performance on these graphs is caused by the different amount of unaccounted noise accumulated in the comparably deeper circuits for these graphs.
@@ IV Discussion
We demonstrate that the post-selected samples with the Iceberg code present a higher approximation ratio than the samples from unencoded circuits. This allows us to study the performance of QAOA in an effective noise regime closer to the noiseless computation; at least for circuits on the beneficial side of the breakeven frontier. We can see the current breakeven frontier of the Iceberg QAOA on 3-regular graphs in Fig. 3A. For example, for $6$ syndrome measurements, the breakeven frontier is up to $k=20$ logical qubits for $\ell\in[4,12]$ of QAOA layers. However, there are multiple opportunities to improve this result and achieve improved performance with Iceberg code as compared to the unencoded circuits for larger qubit counts.
First, other problems may be more ameanable to Iceberg code. Specifically, dealing with sparse-graph problems, like the MaxCut of 3-regular graphs discussed in this paper, is not particularly beneficial for the Iceberg code compared to fully-connected Hamiltonians like the Sherrington-Kirkpatrick model. This is because the Iceberg code can execute two-qubit logical rotations with no overhead, whereas one-qubit logical rotations require a noisier two-qubit physical gate. Iceberg QAOA could be more advantageous for dense graph problems where the number of two-qubit $Z^{\otimes 2}$ terms in the problem Hamiltonian is much larger than the number of single-qubit $X$ terms in the mixing Hamiltonian.
@@
While performance extrapolation indicates promising results with improved hardware fidelity, we observe that, with fixed model parameters, increasing the number of syndrome measurements marginally diminishes the performance gains in the extrapolation heatmaps. Additionally, the overhead of post-selecting samples grows rapidly. This observation is consistent with the experimental results on the H2-1 emulator, indicating that the protective power of the Iceberg error detection code is limited. This reinforces the need for quantum error correction to achieve error rates low enough to run large-scale circuits.
@@ V Methods
This section provides further details on the experimental realization, presents the model construction, and details the fitting to experimental data.
@@ V.1 Location of syndrome measurements
In Fig. 6, we provide evidence supporting this strategy. The $k=16,\ell=11$ logical QAOA circuit is evenly partitioned into eight blocks of logical gates and we introduce a single syndrome measurement in the seven intermediate positions between them. By comparing the logical fidelity obtained from these seven circuits, we find that the circuit with the syndrome measurement inserted in the middle (labeled 4) presents the highest fidelity, at the cost of the smallest post-selection rate.
Figure 6: Top: Circuits with different locations for a single syndrome measurement, labeled 1-7. Bottom: Performance of the different circuits. The circuit with the syndrome measurement in the middle (labeled 4) detects the most errors and achieves the best logical fidelity. All data include the final measurement and fault-tolerant initialization. Without any syndrome measurement, the logical fidelity is $0.5128\pm 0.0115$ and the post-selection rate is $0.2762\pm 0.0058$. With one additional syndrome measurement, all circuits outperform the one without the syndrome measurement.
@@ V.2 Performance model
  • A noise channel with error rate $p_{a}$ that introduces a random error that anti-commutes with both stabilizers after every logical gate of the Iceberg code: insert a random Pauli error from the set $\mathcal{P}^{\otimes 2}\setminus\{I^{\otimes 2},X^{\otimes 2},Y^{\otimes 2},Z^{\otimes 2}\}$ on the gate support.
@@
From the fault-tolerant design every single error in the error detection blocks of the Iceberg code is either detectable or acts trivially, so none of these single errors contribute to the logical infidelity of the Iceberg code. However, two or more errors can produce undetectable errors, that, by virtue of this simplification, produce a logical error. The usually implicit justification of this simplification is that the number of errors that act trivially in the circuit is exponentially smaller in the circuit size than the number of undetectable errors that act non-trivially at the logical level.
@@ V.2.1 Model for the Iceberg code
The model for the Iceberg code incorporates the effects of error detection and fault-tolerance. To construct the model efficiently we divide the circuit into blocks of initialization, logical gates, syndrome measurement and final measurement, respectively, as depicted in Fig. 1. At every block we consider the probabilities of errors causing
@@
    an exciting error $E$ that excites a stabilizer but not an ancilla, propagating such error to the next block without an immediate discard, like a single-qubit $X$ error on a code qubit inserted by the last cnot that acts on that qubit in a syndrome measurement,
@@
In the absence of noise the state remains in the $+1$ eigenspace of the stabilisers and ancillas. An error that excites some of them brings the state to their $-1$ eigenspace, and making the error detectable. We say that exciting and discarding errors are both detectable errors, while harmless and logical errors are undetectable. The sum of the four probabilities adds up to 1 at every block.
@@
These probabilities are initialized from the initialization block, and iteratively updated for every block of logical gates, syndrome measurement, and the final measurement. For example, when adding a block of $g$ logical gates that can suffer internal errors, the input probability of harmless errors $H$ from previous blocks updates to the joint probability $H\leftarrow HP_{0}(p_{c},g)P_{0}(p_{a},g)$ of an input harmless error and no internal errors. At the end of this iterative process we obtain analytical functions for the probabilities of the entire circuit. From those we can compute the model outputs for the Iceberg code.
@@
Errors are evenly distributed across all possible excitation events. That is, for every number $\mu=1$ and $\mu\geq 2$ of errors in a block with $m\geq 0$ ancillas the probability of all excitation events is the same and equal to $\nicefrac{{1}}{{2^{m+2}}}$. The excitation events are the $2^{m+2}$ possible ways to excite (or not) the $m$ ancilla(s) and the two stabilizers.
@@
For example, in the syndrome measurement block there are 16 possible excitation events depending on which of the two stabilizers and the two ancillas are excited or not by errors. Therefore, we assume that $\nicefrac{{1}}{{16}}$ of the errors excite no ancilla and no stabilizer, $\nicefrac{{12}}{{16}}$ of them excite an ancilla, and $\nicefrac{{3}}{{16}}$ excite a stabilizer without exciting the ancillas. This simplification allows to incorporate the error detection and fault-tolerance properties of the Iceberg code in a very natural way. Appendix C show the deviations between the model predictions and the exact fractions computed for small Iceberg code instances.
@@
Initialization block. The top part of Table 2 presents the contribution of every possible error in this block. For each number $\mu=1$ and $\mu\geq 2$ of errors indicated in the first row (columns 2, 3, 4 and columns 3, 4, 5, respectively), the table distinguishes the possible excitations these errors can cause on the ancilla (second row) and stabilizers (third row). Using the simplification 2 we compute the fraction of errors that contribute to each of the possible events and present them in the fourth row. The fifth row shows how this classification incorporates the fault-tolerance of the initialization block. Every single error is either harmless or detectable, while two or more errors that pass undetected (last column) produce a logical error by virtue of simplification 1. Aggregating the contributions provides the initialization of the circuit probabilities presented in the last rows.
@@
Block of logical gates. To add a block of logical gates we need to update the circuit probabilities from the combination of the input errors from the previous blocks with the internal errors of this block. Since no discard is possible at the block of logical gates, the discard probability $D$ is preserved. The top part of Table 3 presents the classification of input, internal anti-commuting and internal commuting errors. The fraction of errors is calculated again using the simplification 2. For undetectable input errors, internal two or more anti-commuting errors cause 4 possible excitation events depending on which of the two stabilizers are excited or not. In contrast, input exciting errors and single anti-commuting errors, each create 3 possible stabilizer excitations: only $S_{Z}$ excited, only $S_{X}$ excited, or both excited. Together they create 9 possible excitation events. In 3 of them the excitations cancel out (11th column) while 6 preserve some excitation (12th column). Similarly, two or more anti-commuting errors create 4 possible excitation events, depending on which of the two stabilizers get excited or not, so here 12 events are possible when combined with an input exciting error. The bottom part of the table presents the resulting update rules of the circuit probabilities to add a block of $g$ logical gates.
@@
Syndrome measurement block. Similarly to the block of logical gates, the contributions from every combination of input and internal errors are summarized in Appendix C. The main difference is that the circuit is immediately discarded if at least one of the two ancillas is excited.
@@
Final measurement block. To add the final measurement block note that the stabilizer $S_{Z}$ is computed in post-process while $S_{X}$ is measured by the ancilla. Note also that the second ancilla of the block is a flag qubit, that can be excited only by internal errors, but not by exciting input errors. Contributions are summarized in Appendix C.
@@
We check in Appendix C that the first and second order terms at low error rates satisfy the expected behaviour
@@
This confirms numerically the partially fault-tolerant nature of the Iceberg code: the only single errors capable of causing a logical error are commuting errors in the logical gates, but no single error from the other two noise sources can cause a logical error.

### v3

@@ III.2.2 Fitted error rates
The deviations between the fitted error rates and the emulator noise rates are presented in Table 1. The error probabilities from the fitted model are larger than the ones from the emulator and provide valuable insights into the accumulation of other noise sources.
Starting from the error rate $p_{cx}$ of cnot gates, the fitted value is more than four times larger than the emulator error rate, showing that a significant amount of noise unaccounted by the performance model accumulates in the error detection blocks of the Iceberg code. For the logical gates, we consider two noise channels with error rates $p_{c}$ and $p_{a}$ that introduce Pauli errors which commute and anti-commute, respectively, with the two Iceberg code stabilizers. Since the values established in the emulator depend on the rotation angles of the logical gates and the QAOA circuits do not present a clear tendency towards any particular angle, Table 1 presents the minimum and maximum values. We find that these fitted error rates are almost double the maximum value given by the emulator, hinting again that the logical gates in the Iceberg code circuit accumulate unaccounted noise. In contrast, the fitted value of the error rate $p_{\ell}$ of the unencoded logical gates is well approximated by the minimum value established by the emulator.
@@
The values reported for the emulator in Table 1 are obtained from the parameters of the depolarising channel for the native two-qubit gates $\exp(-i\theta ZZ)$ and their dependence with the QAOA rotation angles $\theta\in\{\gamma,\beta\}$. The depolarising channel assigns a probability $q_{\sigma}$ to each of the 15 Pauli errors in $\sigma=\mathcal{P}^{\otimes 2}\setminus\{I^{\otimes 2}\}$ after the native gate. The total probability of a Pauli error is the value of $p_{cx}$ reported in the table. These values are corrected by a multiplicative factor defined as a linearly increasing function $r(\theta)\simeq a+b|\theta|$ in the angle magnitude, such that $r(\pi/4)=1$ for maximally entangling gates like the cnot. The error rate of unencoded logical gates is then $p_{\ell}(\theta)=p_{cx}r(\theta)$. Additionally, for the Iceberg code, we separate the error rate of commuting errors $q_{c}=q_{X^{\otimes 2}}+q_{Y^{\otimes 2}}+q_{Z^{\otimes 2}}$ from that of anti-commuting errors $q_{a}=p_{cx}-q_{c}$. To unify with the model error rates, we factorize this single channel into a product of a commuting channel with error rate $p_{c}(\theta)=q_{c}r(\theta)(1-q_{a}r(\theta))$ and an anti-commuting channel with error rate $p_{a}(\theta)=q_{a}r(\theta)$. We report the minimum and maximum among all QAOA rotation angles.
@@
As observed from the shift of the breakeven frontiers (red lines) in Fig. 3A, the QAOA logical fidelity stabilizes as the number of syndrome measurements increases, even though the post-selection rate, indicated by the cyan lines in Fig. 3C, decreases. This aligns with our emulated data and findings in the literature [19], which suggest that the initial syndrome measurements significantly enhance circuit performance, while the marginal gains diminish with an increasing number of syndrome measurements.
Figure 4: Example of using the model to bound the hardware improvement that is necessary but not sufficient for QAOA to become competitive with the Goemans-Williamson algorithm. A Solve $k=16$ MaxCut using different solvers. Each data is reported as the mean of approximation ratio over 100 $k=16$ $3$-regular graphs. The standard errors are too small to be seen. B Scaling of model parameter to beat Goemans-Williamson (GW) algorithm for $k=16$ graphs. The Iceberg code helps the QAOA to beat GW earlier than an unencoded one. The Iceberg and unencoded QAOA data are obtained from the H2-1 emulator.
@@ III.2.4 Predicting performance on future hardware
As shown in Fig. 3B and D, as the factor decreases, we observe a significant shift of the performance frontier to a larger number $k$ of logical qubits while the post-selection rate improves dramatically. This indicates that with higher-quality quantum hardware, we can push the breakeven frontier of logical fidelity to deeper circuits on larger problem instances with less post-selection overhead.
We can use the model to elucidate the necessary but not sufficient conditions for QAOA to become competitive with classical solvers. As an example, we use our model to answer the question of when a QAOA hardware experiment can outperform the Goemans-Williamson (GW) algorithm [60] in terms of approximation ratio on small graphs. As reported in the literature [62], a noiseless QAOA with fixed parameters has been able to surpass the GW algorithm on small graphs. In Fig. 4A, we show an example of solving $k=16$-node 3-regular graphs with noiseless QAOA, GW, and Iceberg QAOA with four syndrome measurements, as well as unencoded QAOA emulated on the H2-1 emulator. The noiseless QAOA is able to outperform GW for $\ell\geq 6$ layers. However, both the Iceberg and unencoded QAOA have not yet surpassed GW. Although more effective heuristic solvers for the MaxCut problem exist [64, 65], we employ the GW algorithm as our baseline due to its robust theoretical foundations and its widespread use in benchmarking quantum algorithms [66, 67, 62]. This discussion does not aim to claim a quantum advantage over classical algorithms. Instead, it illustrates that QED can reduce the physical fidelity requirements needed to match the performance of the GW algorithm.
Specifically, at $\ell=10$, the approximation ratio of noiseless QAOA is $0.9810\ldots$, while the average approximation ratio for GW is $0.9554\ldots$. This implies that the logical fidelity of a noisy QAOA must be of at least $\nicefrac{{0.9554}}{{0.9810}}\simeq 0.974$ to outperform GW, assuming our white noise model approximation. In Fig. 4B, we vary the scaling factor of the model parameters to determine when this breakeven logical fidelity can be achieved. The results indicate that the Iceberg and unencoded circuits require scaling factors of approximately $0.81$ and $0.60$, respectively. Thus, as hardware technology advances, our results suggest that the Iceberg code enables a breakeven fidelity on small MaxCut problems much earlier than unencoded circuits.
Figure 5: Comparison between emulated data and model predictions on random Erdős–Rényi graphs with different numbers of edges. The prediction of Iceberg logical fidelity is less accurate compared to testing on 3-regular graphs, while the prediction of unencoded logical fidelity and the prediction of post-selection rate remain accurate.
@@ III.2.5 Model generalization beyond $3$-regular graphs
We present the comparison between emulated results and model predictions in Fig. 5. The model predictions for the logical fidelity of the Iceberg code circuits are less accurate, specially with the densest graph of $|E|=45$ edges. The model predictions for the unencoded logical fidelity and the Iceberg code post-selection rates are comparably more accurate. This indicates that the fitted model works well for problems with similar topologies, but highlights the limitation of the model’s generalization to different problems. We suspect that the worse model performance on these graphs is caused by the different amount of unaccounted noise accumulated in the comparably deeper circuits for these graphs.
@@ IV Discussion
We demonstrate that the post-selected samples with the Iceberg code present a higher approximation ratio than the samples from unencoded circuits. This allows us to study the performance of QAOA in an effective noise regime closer to the noiseless computation (at least for circuits on the beneficial side of the breakeven frontier). We can see the current breakeven frontier of the Iceberg QAOA on 3-regular graphs in Fig. 3A. For example, for $6$ syndrome measurements, the breakeven frontier is up to $k=20$ logical qubits for $\ell\in[4,12]$ of QAOA layers. However, there are multiple opportunities to improve this result and achieve improved performance with Iceberg code as compared to the unencoded circuits for larger qubit counts.
First, other problems may be more amenable to Iceberg code. Specifically, sparse-graph problems (like the MaxCut of 3-regular graphs discussed in this paper) are not particularly beneficial for the Iceberg code compared to fully-connected Hamiltonians like the Sherrington-Kirkpatrick model (see an example in Appendix C). This is because the Iceberg code can execute two-qubit logical rotations with no overhead, whereas one-qubit logical rotations require a noisier two-qubit physical gate. Iceberg QAOA could be more advantageous for dense graph problems where the number of two-qubit $Z^{\otimes 2}$ terms in the problem Hamiltonian is much larger than the number of single-qubit $X$ terms in the mixing Hamiltonian.
@@
Though performance extrapolation indicates promising results with improved hardware fidelity, we observe that with fixed model parameters, increasing the number of syndrome measurements marginally diminishes the performance gains in the extrapolation heatmaps. Additionally, the overhead of post-selecting samples grows rapidly. This observation is consistent with the emulated results on the H2-1 emulator, indicating that the protective power of the Iceberg error detection code is limited. This reinforces the need for quantum error correction to achieve error rates low enough to run large-scale circuits.
Benchmarking the performance of a quantum algorithm beyond the simulatable region is inherently challenging, particularly when the algorithm is protected by QED or QEC. To address this, we develop a pipeline that calibrates a reliable performance model and extrapolates it to large-scale circuits. We believe this pipeline could inspire further studies on the performance of fault-tolerant implementations of other algorithms.
@@ V Methods
This section provides further details on the circuit realization, presents the model construction, and details the fitting to emulated data.
@@ V.1 Location of syndrome measurements
In Fig. 6, we provide evidence supporting this strategy. The $k=16,\ell=11$ logical QAOA circuit is evenly partitioned into eight blocks of logical gates, and we introduce a single syndrome measurement in the seven intermediate positions between them. By comparing the logical fidelity obtained from these seven circuits, we find that the circuit with the syndrome measurement inserted in the middle (labeled 4) presents the highest fidelity at the cost of the smallest post-selection rate.
Figure 6: Top: Circuits with different locations for a single syndrome measurement, labeled 1-7. Bottom: Emulated performance of the different circuits. The circuit with the syndrome measurement in the middle (labeled 4) detects the most errors and achieves the best logical fidelity. All data include the final measurement and fault-tolerant initialization. Without any syndrome measurement, the logical fidelity is $0.5128\pm 0.0115$ and the post-selection rate is $0.2762\pm 0.0058$. With one additional syndrome measurement, all circuits outperform the one without the syndrome measurement.
@@ V.2 Performance model
  • A noise channel with error rate $p_{a}$ that introduces a random error that anti-commutes with a stabilizer after every logical gate of the Iceberg code: insert a random Pauli error from the set $\mathcal{P}^{\otimes 2}\setminus\{I^{\otimes 2},X^{\otimes 2},Y^{\otimes 2},Z^{\otimes 2}\}$ on the gate support.
@@
From the fault-tolerant design, every single error in the error detection blocks of the Iceberg code is either detectable or acts trivially, so none of these single errors contribute to the logical infidelity of the Iceberg code. However, two or more errors can produce undetectable errors, that, by virtue of this simplification, produce a logical error. The usually implicit justification of this simplification is that the number of errors that act trivially in the circuit is exponentially smaller in the circuit size than the number of undetectable errors that act non-trivially at the logical level.
@@ V.2.1 Model for the Iceberg code
The model for the Iceberg code incorporates the effects of error detection and fault-tolerance. To construct the model efficiently, we divide the circuit into blocks of initialization, logical gates, syndrome measurement and final measurement, respectively, as depicted in Fig. 1. At every block, we consider the probabilities of errors causing
@@
    an exciting error $E$ that excites a stabilizer but not an ancilla, propagating such an error to the next block without an immediate discard, like a single-qubit $X$ error on a code qubit inserted by the last cnot that acts on that qubit in a syndrome measurement,
@@
In the absence of noise, the state remains in the $+1$ eigenspace of the stabilisers and ancillas. An error that excites some of them brings the state to their $-1$ eigenspace, and makes the error detectable. We say that exciting and discarding errors are both detectable errors, while harmless and logical errors are undetectable. The sum of the four probabilities adds up to 1 at every block.
@@
These probabilities are initialized from the initialization block, and iteratively updated for every block of logical gates, syndrome measurement, and the final measurement. For example, when adding a block of $g$ logical gates that can suffer internal errors, the input probability of harmless errors $H$ from previous blocks updates to the joint probability $H\leftarrow HP_{0}(p_{c},g)P_{0}(p_{a},g)$ of an input harmless error and no internal errors. At the end of this iterative process, we obtain analytical functions for the probabilities of the entire circuit, from which we can compute the model outputs for the Iceberg code.
@@
Errors are evenly distributed across all possible excitation events. That is, for every number $\mu=1$ and $\mu\geq 2$ of errors in a block with $m\geq 0$ ancillas, the probability of all excitation events is the same and equal to $\nicefrac{{1}}{{2^{m+2}}}$. The excitation events are the $2^{m+2}$ possible ways to excite (or not) the $m$ ancilla(s) and the two stabilizers.
@@
For example, in the syndrome measurement block, there are 16 possible excitation events, depending on which of the two stabilizers and the two ancillas are excited (or not) by errors. Therefore, we assume that $\nicefrac{{1}}{{16}}$ of the errors excite no ancilla and no stabilizer, $\nicefrac{{12}}{{16}}$ of them excite an ancilla, and $\nicefrac{{3}}{{16}}$ excite a stabilizer without exciting the ancillas. This simplification allows the error detection and fault-tolerance properties of the Iceberg code to be incorporated in a very natural way. Appendix D show the deviations between the model predictions and the exact fractions computed for small Iceberg code instances.
@@
Initialization block. The top part of Table 2 presents the contribution of every possible error in this block. For each number $\mu=1$ and $\mu\geq 2$ of errors indicated in the first row (columns 2, 3, 4 and columns 3, 4, 5, respectively), the table distinguishes the possible excitations these errors can cause on the ancilla (second row) and stabilizers (third row). Using Simplification 2 we compute the fraction of errors that contribute to each of the possible events and present them in the fourth row. The fifth row shows how this classification incorporates the fault-tolerance of the initialization block. Every error is either harmless or detectable, while two or more errors that pass undetected (last column) produce a logical error by virtue of simplification 1. Aggregating the contributions provides the initialization of the circuit probabilities presented in the last rows.
@@
Block of logical gates. To add a block of logical gates, we need to update the circuit probabilities from the combination of the input errors from the previous blocks with the internal errors of this block. Since no discard is possible at the block of logical gates, the discard probability $D$ is preserved. The top part of Table 3 presents the classification of input, internal anti-commuting errors, and internal commuting errors. The fraction of errors is calculated again using Simplification 2. For undetectable input errors, two or more internal anti-commuting errors cause 4 possible excitation events depending on which of the two stabilizers are excited or not. In contrast, input exciting errors and single anti-commuting errors each create 3 possible stabilizer excitations: only $S_{Z}$ excited, only $S_{X}$ excited, or both excited. Together they create 9 possible excitation events. In 3 of them, the excitations cancel out (11th column), while 6 preserve some excitation (12th column). Similarly, two or more anti-commuting errors create 4 possible excitation events, depending on which of the two stabilizers get excited or not. Thus, here 12 events are possible when combined with an input exciting error. The bottom part of the table presents the resulting update rules of the circuit probabilities to add a block of $g$ logical gates.
@@
Syndrome measurement block. Similar to the block of logical gates, the contributions from every combination of input and internal errors are summarized in Appendix D. The main difference is that the circuit is immediately discarded if at least one of the two ancillas is excited.
@@
Final measurement block. To add the final measurement block, note that the stabilizer $S_{Z}$ is computed in post-process, while $S_{X}$ is measured by the ancilla. Note also that the second ancilla of the block is a flag qubit which can be excited only by internal errors (not by exciting input errors). Contributions are summarized in Appendix D.
@@
We check in Appendix D that the first and second order terms at low error rates satisfy the expected behaviour
@@
This numerically confirms the partially fault-tolerant nature of the Iceberg code: the only single errors capable of causing a logical error are commuting errors in the logical gates, but no single error from the other two noise sources can cause a logical error.